solve the proportion 3x 4
-- = --
x + 1 x
: solve the proportion 3x 4
-- = --
x + 1 x

Answer by mangopeeler07(442)

3x/x+1=4/x

Cross multiply:
3x^2=4x+4

Set equal to zero
3x^2-4x-4=0

Factor
(3x+2)(x-2)=0

Solve for x
x=-2/3,2